OpenStudy (anonymous):

Could anyone help with this question: [3*(4)^1/4]-[12*(2)^-3/2] [it would read: 3 times 4 to the (1/4) minus 12 times 2 to the (-3/2)] Sorry about how messy it looks. I have to calculate this using the laws of exponents but I'm not sure where to start.

OpenStudy (anonymous):

the answer is zero: 3*(4)^1/4]-[12*(2)^-3/2 3(4^1/4)=3sqrt(2) & 2^-3/2=12/sqrt(8)=12/2sqrt(2)=6/sqrt2=3sqrt2 and then we have 3sqrt2-3sqrt2=0